OpenStudy (anonymous):

Which explains why the sequence 81, 3, 1/9 is arithmetic or geometric? A) The sequence is arithmetic because it decreases by a factor of 3 B) The sequence is arithmetic because it decreases by a factor of 1/27 C) The sequence is geometric because it decreases by the factor 3 D) The sequence is geometric because it decreases by the factor 1/27

OpenStudy (texaschic101):

arithmetic -- next number is found by adding/subtracting geometric -- next number is found by multiplying to find the next term in a geometric, divide the second term by the first term 3/81 = 0.037 = 1/27 now see if you multiply by 1/27 will it equal the next term 3 * 1/27 = 1/9 3/27 = 1/9 1/9 = 1/9 (correct) and you will see, that to get to the next number, you have to multiply by 1/27 and since you have to multiply to get the next term, it is a geometric sequence with a common ratio of 1/27

OpenStudy (anonymous):

I think I get it. Thanks so much! :D

OpenStudy (texaschic101):

your so welcome :)